OpenStudy (anonymous):

If a:b=2:3, a:c=3:4 and b:d=5:2, find c:d.

OpenStudy (anonymous):

@hartnn

OpenStudy (fifciol):

b=3/2a c=4/3a d=2/5b=3/5a substitute and find value

OpenStudy (anonymous):

how?

OpenStudy (fifciol):

c= 4/3 *a and d=3/5*a so what is c:d?

OpenStudy (fifciol):

\[\frac{ c }{ d }=\frac{ \frac{ 4 }{3 }a}{ \frac{ 3 }{ 5 }a }\]

OpenStudy (anonymous):

20:9 ?

OpenStudy (fifciol):

correct

hartnn (hartnn):

an alternate way : a:b=2:3, a:c=3:4 divide these 2 equations, or simply first flip a:c=3:4 to get the value of c:a then multiply a:b times c:a